OpenStudy (bimbim2697):

WHAT IS THE EQUATION FOR THE LINE THAT IS PARRALEL TO THE GIVEN LINE AND THAT PASSES THROUGH A GIVEN POINT ? Y=1/3X+5/3;(2,1)

OpenStudy (anonymous):

you know m=1/3 So use this formula y-y1=m(x-x1)

OpenStudy (anonymous):

y = 1/3 x + 5/3 In this, slope of the line is 1/3. A line parallel to this line will have same slope. So, the new parallel line is: y = 1/3 x + b (y-intercept is different). Now, to find b, you need to use the information that this new line passes through x=2,y=1 point.

OpenStudy (anonymous):

Since it is parallel, the slope is (1/3) y = (1/3)x + b Since we have the point (2,1), we know that x = 2 when y = 1 1 = (1/3)(2) + b Subtract 2/3 from both sides 1/3 = b Therefore, y = (1/3)x + 1/3

OpenStudy (bimbim2697):

@GT I FOR 3X+5/3

OpenStudy (bimbim2697):

AM I WRONG?

OpenStudy (anonymous):

meverett has the right y-intercept.

OpenStudy (anonymous):

and equation